The cumulative distribution for a random variable x is



F(x) = 1/2 sqrt x-1; 1≤x≤5



Calculate the probabilities



Pr(1≤x≤4.1)

1
Expert's answer
2021-12-02T06:11:33-0500

Given the cumulative distribution,

"F(x)=1/2(\sqrt{x-1})", we can determine the "p(1\leqslant x\leqslant 4.1)" as follows.

"p(1\leqslant x\leqslant4.1)=F(4.1)-F(1)"

Now,

"F(4.1)=(\sqrt{4.1-1})/2=\sqrt{3.1}/2=0.88034084" and "F(1)=(\sqrt{1-1})/2=0"

Therefore,

"p(1\leqslant x\leqslant4.1)=F(4.1)-F(1)=0.88034084-0=0.88034084"

Thus, "pr(1\leqslant x\leqslant 4.1)=0.8803(4dp)"